Reports 1-1 of 1 Clear search Modify search
AdV-INJ (ITF Mode Matching Telescope)
spinicelli, capocasa, rocchi, boldrini, magazzu - 1:11 Wednesday 18 May 2022 (55886) Print this report
Matching of north arm -- estimation of north arm CP focal length

Following the procedure discussed at the commissioning meeting we proceded with the optimization of the north arm matching with the new PR HR power.

We did the following steps:

  1. 1336842894  17 May 2022 17:14:36 UTC        L1z -25.000 (backward)
  2. 1336845574  17 May 2022 17:59:16 UTC        L1z -25.000 (backward)
  3. 1336847533  17 May 2022 18:31:55 UTC        L1z -25.000 (backward)
  4. 1336849469  17 May 2022 19:04:11 UTC        L1z -15.000 (backward)
  5. 1336851939  17 May 2022 19:45:21 UTC        L1z -15.000 (backward)
  6. 1336853947  17 May 2022 20:18:49 UTC        L1z +20.000 (forward)

The evolution of 02 modes and the area below (i.e the mismatching) are shown in pic. 1 and 2 respectively.

Then we switch ON the CH on the west at the power used for compensating the YAG (~150mW).  The plan was to continuolsy scan the modes TEM2 and TEM1 to check both alignment and mismatching, but west green was unstable and unlocked few times (possibly due to a too fast CARM SET scan). Consequently we decided to scan only around TEM2.

A comparison of the modes during the thermal transient is shown in pic 3. 

As we can see, during the thermalization the west arm matching passes through the minimum after more than 15 minutes of CH heating, and then it increases again. We kept scanning the TEM2 mode also during the cooling down after having switched off the WI CH. Data will be further analized offline.

A detailed list of scans and actions performed is in the attached .txt file (see also fig 4)

Images attached to this report
Non-image files attached to this report
Comments to this report:
chiummo - 10:13 Wednesday 18 May 2022 (55889) Print this report

As discussed at last commissioning meeting,  the final tuning of the matching of the input beam to the arms will require the setting of a constant thermally induced lens in the compensation plates. Meanwhile, a check of the matching was performed with no CO2 laser actions.

Started from the situation as left yesterday night, with CH/DAS off. First of all a scan of the arms with green lasers was performed in order to assess the situation.

The matching of the arms looked already pretty good (see fig 1: TEM2 yesterday night, fig 2: TEM2 this morning before changing anything). The meniscus lens on SIB1 was moved by 20000 steps forward which worsen the TEM2(fig 3).

Then the position of the meniscus lens was reversed to explore the other side: moved by 40000 steps backward (fig 4). This improved again the matching comparable to initial value (around 1.5%).

 

To be continued after TCS actions.

Attached a text file with GPS for more accurate evaluation of the mode mismatch, the summary of the actions is reported in fig 5.

 

 

 

 

Images attached to this comment
Non-image files attached to this comment
rocchi - 13:06 Wednesday 18 May 2022 (55895) Print this report

To have the estimate of the North CP cold lens, I made a transient simulation of the evolution of the CH lens (see figure 1).

After 15 minutes, the focal length is about 31 km, already almost equal to the steady state value and half of what was estimated from the CP map.

Images attached to this comment
Search Help
×

Warning

×